A 2-kg sled sliding freely in the positive direction on an icy surface experiences a 4-N frictional force exerted by the ice and an air-resistive force of 0.6 N. Round the final answers to one decimal place.


a. What is the net force acting on the sled?
b. What is the acceleration of the sled?
Answer #1

Given details are: mass of the sled m = 2 kg, frictional force fk = 4 N, air resistive force fr = 0.6 N. Both these forces act opposite to the direction of motion of the sled trying to stop the sled causing negative acceleration i.e deceleration.

a) Net force acting on the sled is Fnet = fk + fr = 4 + 0.6 = 4.6 N, so the resultant force opposing the motion of the sled is Fnet = 4.6 N.

b) From the formula F = ma, the acceleration of the sled is a = F/m = 4.6/2 = 2.3 m/sec2. But since our acceleration is opposite to the motion of the sled, it will be negative so, a = -2.3 m/sec2.